OpenStudy (anonymous):

help (: determine y in log base 16, if y=1/2

OpenStudy (anonymous):

a little confusing because you say "determine y" yet there is \[y=\frac{1}{2}\]

OpenStudy (anonymous):

unless you mean \[\log_{16}(y)=\frac{1}{2}\]?

OpenStudy (anonymous):

yeah the 2nd part lol

OpenStudy (anonymous):

in which case you write this in equivalent exponential form which is \[16^{\frac{1}{2}}=y\]

OpenStudy (anonymous):

and \[16^{\frac{1}{2}}=\sqrt{16}=4\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

i worked it out and got "y" = 4

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

\[log_b(y)=x\] same as \[b^x=y\]

OpenStudy (anonymous):

yeah that is what i got too!

OpenStudy (anonymous):

we're genuises (: